# Human-in-the-Loop Causal Discovery under Latent Confounding using Ancestral GFlowNets

### v2

@@ Appendix B Cross-entropy acquisition
$\begin{split}\mathbf{H}&[q(\cdot|\mathbf{f}_{K+1}),q(\cdot|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}[-\log q(\mathcal{G}|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[-\log p_{\theta}(\mathcal{G})-\sum_{1\leq k\leq K}\log p(\omega_{r_{k}}|f_{r_{k}})-l_{K}\right]\end{split},$
@@
in which $l_{K}$ is the log-partition function of the distribution $q(\cdot|\mathbf{f}_{K})$. Further, the cross-entropy depends exclusively upon i) the logarithm of the samples’ rewards, $\log p_{\theta}(\mathcal{G})$, which is readily computed within AGFN’s generative process, and ii) the posterior distribution over the relations’ features $\omega_{r}$ given the expert’s feedbacks $f_{r}$, which is available in closed form. Hence, the previously mentioned expectation is unbiasedly and consistently estimated by our importance sampling scheme. Furthermore, our empirical findings in fig. 5 suggest that the cross-entropy yields good results and consistently outperforms a uniformly random strategy with respect to the BIC score.
@@ Appendix C Experimental details
We lay out the experimental and implementational details of our empirical analysis in the next subsections. In Section C.1, we describe the specific configurations of the CD algorithms that we compared with our method in table 2. Then, we consider in section C.2 some practical guidelines and architectural specifications that enable us to train and make inferences with AGFN efficiently. Finally, we contemplate in section C.3 the algorithmic details for simulating the expert’s feedback according to our model for active knowledge elicitation.
@@ C.1 Baselines
For the results in table 1, we first estimated a PAG using the stable version of FCI, which produces a fully order-independent final skeleton (Colombo et al., 2014). To identify conditional independencies, we used Fisher’s Z partial correlation test with a significance level of $\alpha=0.05$. The BIC score associated with the PAG estimated by the FCI was computed as the BIC of a randomly selected maximal AG (MAG) within the equivalence class characterized by such PAG. The maximality of an AG depends on the absence of inducing paths between non-adjacent variables, which are paths where every node along it (except the endpoints) is a collider and every collider is an ancestor of an endpoint (Rantanen et al., 2021). This ensures that in the MAG every non-adjacent pair of nodes is m-separated by some set of other variables. Importantly, Markov equivalent MAGs exhibit asymptotic equivalence in terms of BIC scores (Richardson and Spirtes, 2002). As a result, the choice of a random MAG does not disrupt the validity of our results.
GFCI.
@@
Similarly, we applied GFCI with an initial search algorithm (FGS) based on the BIC score and the subsequent application of the FCI with conditional independencies identified by the Fisher’s Z partial correlation test with a significance level $\alpha=0.05$. This was performed for all datasets listed in table 1. Also similar to the procedure adopted with the FCI, the BIC score associated with the estimated PAG was computed as the BIC of a randomly selected MAG within the equivalence class characterized by such PAG.
DCD.
@@
We adhered to the instructions provided in the official repository [^1] Available online at https://gitlab.com/rbhatta8/dcd. to apply the DCD method on the datasets in table 1. The SHD was obtained between the ground-truth PAG and the PAG corresponding to the estimated ADMG (i.e., the one obtained via FCI by using the d-separations entailed by the estimated ADMG as an oracle for conditional independencies). On the other hand, the BIC was computed for the estimated ADMG directly.
N-ADMG.
@@
To estimate the parameters of the variational distribution defined by N-ADMG, we executed the code provided at the official repository [^2] Available online at https://github.com/microsoft/causica/releases/tag/v0.0.0. For fairness, we used the same hyperparameters and architectures reported in their original work Li et al. (2023); in particular, we trained the models for $30$k epochs. After this, we sampled $100$k graphs from the learned distribution. It is worth mentioning that the constraints of bow-free ADMG are guaranteed in the N-ADMG samples only in an asymptotic sense. Thus, we manually removed any cyclic graphs from the learned distribution. Then, we proceeded exactly as with DCD to estimate both the average SHD and the average BIC under the variational distribution.
C.2 Implementational details for AGFN
@@
Masking.
To ensure AGFN only samples ancestral graphs, we keep track of a binary mask $\mathbf{m}_{t}$ that indicates which actions lead to a valid state at the iteration $t$ of the generative process; this mask defines the support of the policy evaluated at the corresponding state. In more detail, let $\mathbf{y}_{t}$ be the last layer embedding (prior to a softmax) at iteration $t$ of the neural network used to parametrize the forward flow of AGFN. The probability distribution over the space of feasible actions is then
@@
for a large and negative constant $\epsilon$. We empirically verified that $\epsilon=-10^{5}$ is sufficient to avoid the sampling of non-ancestral graphs.
Exploratory policy.
During training, we must use an exploratory policy that (i) enables the exploration of yet unvisited states within the pointed DAG and (ii) exploits highly valuable and already visited states. To alleviate this phenomenon, we also draw trajectories from a uniform policy, which is a widespread practice in the literature (Bengio et al., 2021a; Deleu et al., 2022; Shen et al., 2023). More precisely, let $\text{Ch}(\mathcal{G}_{t})$ be the set of states (i.e., ancestral graphs) directly reachable from $\mathcal{G}_{t}$ and $\alpha\in[0,1]$. At each iteration $t$ of the generative process, we sample an action (either an edge to be appended to the graph or a signal to stop the process)
@@
Detection of invalid states.
We use the algebraic condition in eq. 8 to check whether a graph $\mathcal{G}$ is ancestral. At each iteration of the generative process, we draw an action from the current exploratory policy and test the ancestrality of the updated graph; if it is not ancestral, we revert the sampled action and mask it. Importantly, this protocol guarantees that all graphs sampled from AGFN are ancestral.
Batch sampling.
@@
We exploit batch sampling to fully leverage the power of GPU-based computing in AGFN. As both the maximum-log-likelihood-based reward and the validation of the states are parallelizable operations, we are able to distribute them across multiple processing units and efficiently draw samples from the learned distribution. Crucially, this end-to-end parallelization substantially improves the computational feasibility of our algorithm and is a notable feature generally unavailable in prior works Zhang (2008b); Ogarrio et al. (2016); Rantanen et al. (2021). We use a batch size of 256 for all the experiments — independently of the graph size.
@@
Figure 7: Sensitivity of our active knowledge elicitation framework to the reliability of the expert. Each column represents either the expected SHD (top) or expected BIC (bottom) as a function of the degree of confidence $\pi\in[0,1]$ in the expert as a function of the number of feedbacks. As expected, the improvements entailed by the expert’s feedback become increasingly effective as we increase the expert’s reliability from $0.1$ to $0.9$. Results reflect the outcome of 30 scenarios simulated accordingly to algorithm 1 with a random canonical diagram $\mathcal{G}^{\star}$ with $5$ nodes. We used our active knowledge elicitation scheme to select the query at each iteration.
Training hyperparameters.
@@
For AGFN’s forward flow, we use a Graph Isomorphism Network (GIN, Xu et al. (2019)) with $2$ layers to compute embeddings of dimension $256$. Then, we project these embeddings to a probability distribution using a three-layer MLP having leaky RELUs with a negative slope of $-0.01$ as activation functions. Correspondingly, we use an equally configured three-layer MLP to parametrize AGFN’s backward flow. For training, we use the Adam method for the stochastic optimization problem defined by the minimization of the loss in eq. 7. Moreover, we trained the neural networks for $3000$ epochs for the human-in-the-loop simulations (in which we considered graphs having up to $10$ nodes) and for $500$ epochs for both the assessment of the distributional quality of AGFN and the comparison of AGFN with alternative CD approaches.
Computational settings.
@@
We trained the AGFNs for the experiments in fig. 3 and table 1 and fig. 5 for $500$ epochs in computers equipped with NVIDIA’s V100 GPUs. All the experiments were executed in a cluster of NVIDIA’s V100 GPUs and the algorithms were implemented using the machine learning framework PyTorch. To estimate the PAG corresponding to AGFN’s samples and compute the SHDs reported in table 1, we used the FCI’s implementation of the pcalg package in R considering the d-separations entailed by these samples as a criterion for conditional dependence.
C.3 Human in the loop
Algorithm 1 Simulating humans in the loop
$\{\mathcal{G}_{t}\}_{1\leq t\leq T}\text{ samples from AGFN}$, $\mathcal{G}^{*}=(\mathbf{V},E)$ true ancestral graph, $\pi$ reliability of the expert’s feedback
$p(\omega_{r}=k)\leftarrow\frac{1}{T}\sum_{1\leq t\leq T}1_{\{\omega_{r}=k\text{ in }G_{t}\}}\forall k\in[4],r\in{\mathbf{V}\choose 2}$
$\mathbf{f}\leftarrow\{\}$triangleright Set of feedbacks (answers)
$\boldsymbol{r}\leftarrow\{\}$triangleright Set of queries (questions)
$K\leftarrow 1$
while $\boldsymbol{r}\neq{\mathbf{V}\choose 2}$ do $\triangleright$ Iteratively request the feedback
     $\mathbf{r}\leftarrow\mathbf{r}\cup\{r_{K}\}$
     $\mathbf{f}\leftarrow\mathbf{f}\cup\{f_{K}\}$
     $K\leftarrow K+1$
end while
Algorithmic details.
We describe in algorithm 1 our procedure for simulating interactions with an expert. Initially, we estimate the marginal probabilities $p(\omega_{r}=k)$ of a relation $r$ displaying the feature $k\in\{1,2,3,4\}$ under AGFN’s learned distribution. This is our prior distribution. In algorithm 1, we denote $\{1,2,3,4\}$ by $[4]$. Then, we iteratively select the relation that maximizes our acquisition function; the simulated human thus returns a feedback that equals the selected relation’s true feature with probability $\pi$ or is otherwise uniformly distributed among the incorrect alternatives. Importantly, this iterative mechanism can be interrupted at any iteration and the collected feedbacks can be used to compute the importance weights necessary for estimating expectations of functionals under our updated beliefs.
Appendix D Additional experiments

### v3

@@ A.3 Cross-entropy
$\begin{split}\mathbf{H}&[q(\cdot|\mathbf{f}_{K+1}),q(\cdot|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}[-\log q(\mathcal{G}|\mathbf{f}_{K})]\\ &=\underset{\mathcal{G}\sim q(\cdot|\mathbf{f}_{K+1})}{\mathbb{E}}\left[-\log p_{\theta}(\mathcal{G})-\sum_{1\leq k\leq K}\log p(\omega_{r_{k}}|f_{r_{k}})-l_{K}\right]\end{split},$ (21)
@@
in which $l_{K}$ is the log-partition function of the distribution $q(\cdot|\mathbf{f}_{K})$, there is no need to estimate any partition function for selecting a query with which to probe the expert. Further, the cross-entropy depends exclusively upon i) the logarithm of the samples’ rewards, $\log p_{\theta}(\mathcal{G})$, which is readily computed within AGFN’s generative process, and ii) the posterior distribution over the relations’ features $\omega_{r}$ given the expert’s feedbacks $f_{r}$, which is available in closed form. Hence, the previously mentioned expectation in Equation 21 is unbiasedly and consistently estimated through a Monte Carlo estimator based on samples drawn from AGFN’s updated distribution. Furthermore, our empirical findings in fig. 12 suggest that the cross-entropy yields good results and consistently outperforms a uniformly random strategy with respect to the BIC score.
Appendix B Algorithmic design of AGFN
To keep track of the invalid actions leading to non-ancestral graphs within the generative process, we follow an approach similar to Giudici and Castelo, (2003); Deleu et al., (2022) and decompose the mask $\mathbf{m}_{t}$ of valid actions at the $t$th state $G_{t}$ into a term corresponding to the state’s adjacency matrix and a term resembling the state’s transitive closure. Formally, we define
  1. 1.
    $\alpha_{t}\colon\mathbf{V}\times\mathbf{V}\rightarrow\{1,0\}$ as an indicator function verifying whether there is a directed path between the nodes $v_{i}\in V$ and $v_{j}\in\mathbf{V}$ in the transpose graph $\mathcal{G}_{t}^{\intercal}$; and
  2. 2.
    $\beta_{t}\colon\mathbf{V}\times\mathbf{V}\rightarrow\{1,0\}$ as another indicator function identifying whether there is, within the transpose graph, a path between $v_{i}\in\mathbf{V}$ and $v_{j}\in\mathbf{V}$ with all, except one, directed edges pointing to $v_{j}$.
In this context, we note that adding a directed edge $v_{i}\rightarrow v_{j}$ leads to an cycle if and only if $\alpha_{t}(v_{i},v_{j})=1$, whereas it leads to an almost cycle if and only if $\beta_{t}(v_{i},v_{j})=1$. Correlatively, the addition of a bidirected edge $v_{i}\leftrightarrow v_{j}$ induces an almost cycle within the current state if and only if either $\alpha_{t}(v_{i},v_{j})=1$ or $\alpha_{t}(v_{j},v_{i})=1$. In this context, as an action is allowed if and only if it doesn’t lead to the formation of a directed or an almost directed cycle, we split the mask $\mathbf{m}_{t}=\mathbf{d}_{t}\oplus\mathbf{b}_{t}$ as a concatenation into actions corresponding to adding directed edges $\mathbf{d}_{t}$ and actions corresponding to bidirected edges $\mathbf{b}_{t}$, decomposing it as
$\begin{split}&\mathbf{d}_{t}=\mathbf{E}_{t}\odot\left(1-\boldsymbol{\alpha}_{t}\right)\odot\left(1-\boldsymbol{\beta}_{t}\right)\\ &\mathbf{b}_{t}=\mathrm{triu}\left(\mathbf{E}_{t}\odot\left(1-\boldsymbol{\alpha}_{t}\right)\odot\left(1-\boldsymbol{\alpha}_{t}^{\intercal}\right)\right)\end{split},$ (22)
with $\mathbf{E}_{t}=\left(1-\mathbf{D}_{t}\right)\odot\left(1-\mathbf{B}_{t}\right)$ denoting the unconnectedness of each pair of nodes; $\boldsymbol{\alpha}_{t}$ (resp. $\boldsymbol{\beta}_{t})$) representing a matrix with entries $\alpha_{t}(v_{i},v_{j})$ (resp. $\beta_{t}(v_{i},v_{j})$); and $\mathrm{triu}$ corresponding to the operation of extracting the upper triangular part of a matrix. Notably, each of the terms of the preceding equation can be efficiently updated as we iteratively build the ancestral graphs. For $\mathbf{E}_{t}$, we simply update the entries corresponding to the chosen action. For $\boldsymbol{\alpha}_{t}$, notice that adding an edge $v_{i}\rightarrow v_{j}$ creates a directed path from the node $w_{j}$ to the node $w_{i}$ in $\mathcal{G}_{t}^{\intercal}$ if and only if there is a directed path from $w_{j}$ to $v_{j}$ and a directed path from $v_{i}$ to $w_{i}$ in $\mathcal{G}_{t}^{\intercal}$, with a similar correspondence for directed paths with single bidirected edges; thus, after adding $w_{i}\rightarrow w_{j}$, we update
$\begin{split}&\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\boldsymbol{\alpha}_{t}(\cdot,v_{j})\boldsymbol{\alpha}_{t}(v_{i},\cdot)\\ &\boldsymbol{\beta}_{t+1}=7\boldsymbol{\beta}_{t}+\boldsymbol{\beta}_{t}(\cdot,v_{j})\boldsymbol{\alpha}_{t}(v_{i},\cdot)+\boldsymbol{\alpha}_{t}(\cdot,v_{j})\boldsymbol{\beta}_{t}(v_{i},\cdot)\end{split},$ (23)
in which $\boldsymbol{\alpha}_{t}(\cdot,v)$ (resp. $\boldsymbol{\alpha}_{t}(v,\cdot)$) corresponds to the column (resp. row) vector indexed by $v$ in $\boldsymbol{\alpha}_{t}$ and the operator ++ represents a bitwise or. By a similar reasoning, the addition of a bidirected edge $v_{i}\leftrightarrow v_{j}$ corresponds to the update rule
$\begin{split}&\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}\\ &\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\boldsymbol{\alpha}_{t}(\cdot,v_{i})\boldsymbol{\alpha}_{t}(v_{j},\cdot)+\boldsymbol{\alpha}_{t}(\cdot,v_{j})\boldsymbol{\alpha}_{t}(v_{i},\cdot)\end{split}.$ (24)
Alternatively, albeit significantly more costly, one could enumerate the potential children of a state and use Bhattacharya et al., (2021)’s algebraic constraint to decide the validity of the corresponding transitions.
@@ Appendix C Experimental details
We lay out more experimental and implementation details of our empirical analysis in the next subsections.
@@ C.1.1 FCI
We first estimated a PAG using the stable version of FCI, which produces a fully order-independent final skeleton (Colombo et al.,, 2014). To identify conditional independencies, we used Fisher’s Z partial correlation test with a significance level of $\alpha=0.05$. The BIC score associated with the PAG estimated by the FCI was computed as the BIC of a randomly selected maximal AG (MAG) within the equivalence class characterized by such PAG. The maximality of an AG depends on the absence of inducing paths between non-adjacent variables, which are paths where every node along it (except the endpoints) is a collider and every collider is an ancestor of an endpoint (Rantanen et al.,, 2021). This ensures that in the MAG every non-adjacent pair of nodes is m-separated by some set of other variables. Importantly, Markov equivalent MAGs exhibit asymptotic equivalence in terms of BIC scores (Richardson and Spirtes,, 2002). As a result, the choice of a random MAG does not disrupt the validity of our results.
C.1.2 GFCI
@@
Similarly, we applied GFCI with an initial search algorithm (FGS) based on the BIC score and the subsequent application of the FCI with conditional independencies identified by the Fisher’s Z partial correlation test with a significance level $\alpha=0.05$. Also similar to the procedure adopted with the FCI, the BIC score associated with the estimated PAG was computed as the BIC of a randomly selected MAG within the equivalence class characterized by such PAG.
C.1.3 ACI
We used the implementation provided by the authors at GitHub [^1] ACI repository, available online at https://github.com/caus-am/aci. For the Answer Set Programming solver, we used clingo 4, version 5.6.2, also available at GitHub  [^2] Clingo repository, available online at https://github.com/potassco/clingo. Similarly to the FCI, we tested conditional indepedencies using Fisher’s Z partial correlation test with a significance level of $\alpha=0.05$. ACI also outputs a PAG, so we computed the BIC of a randomly selected MAG within the represented Markov equivalence class.
C.1.4 DCD
@@
We adhered to the instructions provided in the official repository [^3] Available online at https://gitlab.com/rbhatta8/dcd. to apply the DCD method. The SHD was obtained between the ground-truth PAG and the PAG corresponding to the estimated ADMG (i.e., the one obtained via FCI by using the d-separations entailed by the estimated ADMG as an oracle for conditional independencies). On the other hand, the BIC was computed for the estimated ADMG directly.
C.1.5 N-ADMG
@@
To estimate the parameters of the variational distribution defined by N-ADMG, we executed the code provided at the official repository [^4] Available online at https://github.com/microsoft/causica/releases/tag/v0.0.0. For fairness, we used the same hyperparameters and architectures reported in their original work (Li et al.,, 2023); in particular, we trained the models for 3030k epochs. After this, we sampled 100100k graphs from the learned distribution. It is worth mentioning that the constraints of bow-free ADMG are guaranteed in the N-ADMG samples only in an asymptotic sense. Thus, we manually removed any cyclic graphs from the learned distribution. Then, we proceeded exactly as with DCD to estimate both the average SHD and the average BIC under the variational distribution.
C.1.6 DAG-GFlowNets
DAG-GFlowNets (Deleu et al.,, 2022) are a GFlowNet-based method for structure learning of Bayesian networks. In a nutshell, this family of models first defines a generative process of DAGs by iteratively adding edges to an initially edgeless graph. Then, a GFlowNet is trained to sample states proportionally to a goodness-of-fit score function (Deleu et al.,, 2023). Importantly, when regarded as a CD algorithm, DAG-GFlowNets ignore latent confounding. To illustrate this limitation, we consider three distinct causal graphs belonging to MEC with invariant bidirected edges (see diagrams 1, 2, and 3 in Figure 6). Under these conditions, we show that AGFNs are often more accurate than DAG-GFlowNets.
Results To evaluate the score function in Equation 27, we considered datasets with 500 samples. Table 2 below compares the SHD of point estimates for DAG-GFlowNets with AGFN’s. Notably, AGFN outperforms DAG-GFlowNet in all cases. This is expected, since DAG-GFlowNets, in contrast with AGFNs, are incapable of sampling/learning graphs with bidirected edges, even under faithfulness.
DAG-GFlowNet AGFN
Causal diagram 1 $3.80$ $\mathbf{1.67}$
Causal diagram 2 $6.40$ $\mathbf{5.27}$
Causal diagram 3 $8.22$ $\mathbf{7.87}$
Table 2: AGFN outperforms DAG-GFN under the presence of latent confounders. See Figure 6 for the diagrams.
(a) CD 1.
(b) CD 2.
(c) CD 3.
Figure 6: Causal diagrams under latent confounding for comparing AGFN with DAG-GFlowNet. See results in Table 2.
C.2 Implementation details for AGFN
C.2.1 Forward flow
We use a Graph Isomorphism Network (Xu et al.,, 2019) $\Phi$ to compute a $d$-dimensional representation for each node in the AG $\mathcal{G}_{t}$ at the $t$-th step of the generative process and use sum pooling to get an embedding for $\mathcal{G}_{t}$. Let $\mathcal{A}_{t}$ denote the space of feasible actions at $\mathcal{G}_{t}$ (i.e., those leading to valid AGs). We then map $\mathcal{G}_{t}$’s embedding to a distribution over $\mathcal{A}_{t}$ using a multilayer perceptron (MLP) $\phi\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{A}_{t}|}$ with a softmax activation at its last layer. Specifically, given $\mathbf{H}^{(t)}=\Phi(\mathcal{G}_{t})\in\mathbb{R}^{|\mathbf{V}|\times d}$, we compute
$\mathbf{p}=\phi\left(\sum\nolimits_{v\in\mathbf{V}}\mathbf{H}_{v}^{(t)}\right)$ (25)
as the probability distribution over the actions at $\mathcal{G}_{t}$.
C.2.2 Backward flow
Backward actions correspond to removing edges. Following Shen et al., (2023), we parametrize the backward flow $F_{B}$ with an MLP and alternate between updating $\pi_{F}$ and $\pi_{F_{B}}$, using gradient-based optimization.
@@ C.2.2 Backward flow
C.2.3 Masking
To ensure AGFN only samples ancestral graphs, we keep track of a binary mask $\mathbf{m}_{t}$ that indicates which actions lead to a valid state at the iteration $t$ of the generative process; this mask defines the support of the policy evaluated at the corresponding state. See Appendix B for further details regarding the computation of this mask. In practice, we let $\mathbf{y}_{t}$ be the last layer embedding (prior to a softmax) at iteration $t$ of the neural network used to parametrize the forward policy of AGFN. The probability distribution over the space of feasible actions is then
@@ C.2.3 Masking
for a large and negative constant $\epsilon$. We empirically verified that $\epsilon=-10^{5}$ is sufficient to avoid the sampling of non-ancestral graphs when using a double-precision floating point format.
C.2.4 Exploratory policy
During training, we must use an exploratory policy that (i) enables the exploration of yet unvisited states within the pointed DAG and (ii) exploits highly valuable and already visited states. To satisfy these criteria, we also draw trajectories from a uniform policy, which is a widespread practice in the literature (Bengio et al., 2021a, ; Deleu et al.,, 2022; Shen et al.,, 2023). More precisely, let $\text{Ch}(\mathcal{G}_{t})$ be the set of states (i.e., ancestral graphs) directly reachable from $\mathcal{G}_{t}$ and $\alpha\in[0,1]$. At each iteration $t$ of the generative process, we sample an action (either an edge to be appended to the graph or a signal to stop the process)
@@ C.2.4 Exploratory policy
C.2.5 Validating AGFN’s samples
To validate that AGFN’s sampled graphs are ancestral, we use the algebraic condition proposed by Bhattacharya et al., (2021), namely,
$\textrm{trace}(e^{D})-d+\mathbf{1}^{T}e^{D\odot B}\cdot\mathbf{1}=0,$ (26)
in which $D\in\{1,0\}^{d}$ is the graph’s adjacency matrix corresponding to the directed edges and $B$ is the one corresponding to the bidirected edges; $\mathbf{1}$ is a vector of ones; and $\odot$ represents Hadamard’s pointwise product.
C.2.6 Batch sampling
@@
We exploit batch sampling to fully leverage the power of GPU-based computing in AGFN. As both the maximum-log-likelihood-based reward and the validation of the states are parallelizable operations, we are able to distribute them across multiple processing units and efficiently draw samples from the learned distribution. Crucially, this end-to-end parallelization substantially improves the computational feasibility of our algorithm and is a notable feature generally unavailable in prior works (Zhang, 2008b, ; Ogarrio et al.,, 2016; Rantanen et al.,, 2021). We use a batch size of 128 trajectories for all the experiments, independently of the graph size.
@@
Figure 8: Sensitivity of our active knowledge elicitation framework to the reliability of the expert. Each column represents either the expected SHD (top) or expected BIC (bottom) as a function of the degree of confidence $\pi\in[0,1]$ in the expert as a function of the number of feedbacks. As expected, the improvements entailed by the expert’s feedback become increasingly effective as we increase the expert’s reliability from $0.1$ to $0.9$. Results reflect the outcome of 30 scenarios simulated accordingly to algorithm 1 with a random canonical diagram $\mathcal{G}^{\star}$ with 55 nodes. We used our active knowledge elicitation scheme to select the query at each iteration.
C.2.7 Training hyperparameters
@@
For AGFN’s forward flow, we use a Graph Isomorphism Network (GIN, Xu et al., (2019)) with $2$ layers to compute embeddings of dimension $256$. Then, we project these embeddings to a probability distribution using a three-layer MLP having leaky RELUs with a negative slope of 0.01-0.01 as activation functions. Correspondingly, we use an equally configured three-layer MLP to parametrize AGFN’s backward flow. For training, we use the Adam method for the stochastic optimization problem defined by the minimization of the loss in eq. 7. Moreover, we trained the neural networks for $3000$ epochs for the human-in-the-loop simulations (in which we considered graphs having up to $10$ nodes) and for $500$ epochs for both the assessment of the distributional quality of AGFN and the comparison of AGFN with alternative CD approaches.
C.2.8 Computational settings
@@
We trained the AGFNs for the experiments for $500$ epochs in computers equipped with NVIDIA’s V100 GPUs. All the experiments were executed in a cluster of NVIDIA’s V100 GPUs and the algorithms were implemented using the machine learning framework PyTorch (Paszke et al.,, 2019). To estimate the PAG corresponding to AGFN’s samples and compute the SHDs, we used the FCI’s implementation of the pcalg package in R considering the d-separations entailed by these samples as a criterion for conditional dependence.
C.2.9 Running times for AGFN
Figure 9 highlights that the required training and sampling times for AGFN scale roughly as a quadratic function of the number of observed variables. Correspondingly, the running time for drawing independent graphs from the learned distribution increases linearly in the number of graphs. Contrastingly, the cost of SOTA CD algorithms grows exponentially in the number of variables (Zhang, 2008b, ; Ogarrio et al.,, 2016), and sampling from the bootstrapped distribution is thus prohibitively expensive.
Figure 9: Running times for AGFN. The avg. time AGFN takes to train/sample scales $\approx$ quadratically in the number of observed variables ($\leftarrow$, center). Similarly, the sampling time grows linearly with the number of sampled graphs ($\rightarrow$).
C.2.10 Score for linear Gaussian models
For experiments with linear Gaussian models, we use the extended Bayesian Information Criterion (Foygel and Drton,, 2010) as our score function. Specifically, for any AG $\mathcal{G}=(\mathbf{V},\mathbf{E})$:
[…]
C.3 Sachs dataset details
Sachs (Sachs et al.,, 2005) is a dataset that measures the levels of specific proteins and phospholipids in the human cell. The data are continuous and the $11$ observed variables/features are: "Raf", "Mek", "Plcg", "PIP2", "PIP3", "Erk", "Akt", "PKA", "PKC", "P38", "Jnk".
As we would like the dataset to contain latent counfounders, we used three different versions in our experiments: Sachs, Sachs-7 and Sachs-5. All have an equal number of samples, $7466$, which is the original amount.
Sachs is the original dataset. In order to reduce the number of variables, we apply a simple conditional independence test and remove the ones with the most false independencies. Thus, we ensure the resulting graphs remain sufficiently connected and not overly sparse.
Sachs-7 contains the following variables: "Mek", "Plcg", "PIP2", "PIP3", "PKA", "PKC", "Jnk" and Sachs-5: "Mek", "PIP3", "PKA", "PKC", "Jnk".